[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed May 17, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed May 17 16:47:56 UTC 2017


HSU Shift Log for Wed May 17 2017 12:47:50 GMT-0400 (EDT)

Shift: H

Forecaster: M Brennan

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: I noticed that the 48-h EP consensus graphic on the  
FSU genesis website hadn't updated since 11 May, so I emailed Bob Hart and  
Dan Halperin about it. http://moe.met.fsu.edu/modelgen/con48epac.php The  
5-day graphic seems OK however.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170517/c82a3b85/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list